A Bibliography of Publications of Jean Louchet

Jean Louchet ENSTA LEI 32 boul Victor 75739 Paris 15 France

Tel: +33-1-4552 6075 FAX: +33-1-4552 5587

E-mail: Jean.Louchet@inria.fr (Internet)

25 November 2011 Version 1.02

Abstract

This bibliography records publications of Jean Louchet.

Title word cross-reference

algorithm [Lou94]. analysis [Lou94]. animation [LBCP95, LPC95].

Building [LBCP95].

cloth [LPC95].

Evolutionary [LPC95, Lou94, LBCP95].

identification [LPC95]. image [LBCP95].

models [LPC95]. motion [Lou94].

physical [Lou94]. synthetic [LBCP95].

techniques [LBCP95]. tools [LBCP95].

using [LBCP95].

References

Louchet:1995:BNT

[LBCP95] J. Louchet, M. Boccara, D. Crochemore, and X. Provot. Building new tools for synthetic image animation using evolutionary techniques. Artificial Evolution ’95 (Springer-Verlag), to appear), September 1995.

Louchet:1994:EAP

[Lou94] J. Louchet. An evolutionary algorithm for physical motion analysis.

British Machine Vision Conference, 2:701–710, September 1994.

Louchet:1995:EIC

[LPC95] J. Louchet, X. Provot, and D. Crochemore. Evolutionary identification of cloth animation models. Computer Animation and Simulation ’95 (Springer-Verlag), pages 44–54, September 1995.